OpenStudy (anonymous):

When 1 volume of gas X reacts with exactly 5 volumes of oxygen it forms carbon dioxide and water only. What is gas X? A methane, CH4 B ethane, C2H6 C propane, C3H8 D butane, C4H10

OpenStudy (anonymous):

Please show me an easy way to do this!

OpenStudy (anonymous):

it is C propane to find the number of moles of Oxygen there is a general formula CxHy X+Y/2 = Number of Oxygen required

OpenStudy (anonymous):

I didn't get that formula. @hashsam1

OpenStudy (anonymous):

you know co2 and h20 are formed when any hydrocarbon undergoes Combustion so if want to burn any hydrocarbon CxHy +O2= CO2 + H20 TO BALANCE IT O2= X+Y/4 CO2=X H20=Y/2

OpenStudy (anonymous):

Oh. Ok. Thank you so much! :)

OpenStudy (anonymous):

Np